in a certain factory there are 10 officers, 60 clerks and 400 workers.Their average salary are 627,230 and 120 respectively. Find the average salary of the employees working together
1
Expert's answer
2019-11-26T09:52:05-0500

The average salary of 10 officers is 627. So the total salary of 10 0fficers will be "(10*627)=6270"

The average salary of 60 clerks is 230, so their total salary will be "(60*230)=13800"

The average salary of 400 workers is 120, so their total salary will be "(400*120)=48000"

Now, the total salary of all the employees will be "(6270+13800+48000)=68070"

And the total number of employees altogether is "(10+60+400)=470"

So, the average salary of all the employee working together will be "(68070/470)\approx144.83"